# 1]Fundamental AI Research (FAIR) at Meta

### v1

@@ Abstract
Machine learning interatomic potentials (MLIPs) have become increasingly effective at approximating quantum mechanical calculations at a fraction of the computational cost. However, lower errors on held out test sets do not always translate to improved results on downstream physical property prediction tasks. In this paper, we propose testing MLIPs on their practical ability to conserve energy during molecular dynamic simulations. If passed, improved correlations are found between test errors and their performance on physical property prediction tasks. We identify choices which may lead to models failing this test, and use these observations to improve upon highly-expressive models. The resulting model, eSEN, provides state-of-the-art results on a range of physical property prediction tasks, including materials stability prediction, thermal conductivity prediction, and phonon calculations.
@@ 1 Introduction
Figure 1: (a) Energy conservation in MD simulations. Direct-force models (Orb, eqV2) and CHGNet fail to conserve. (b) A higher F-1 score on the Matbench-Discovery strongly correlates with a lower test-set energy MAE. (c) Test-set energy MAE and $\kappa_{\mathrm{SRME}}$ on the Matbench-Discovery benchmark. (d) Test-set energy MAE and vibrational entropy MAE on the MDR Phonon benchmark. Our model (eSEN) achieves the best performance on all benchmarks. A higher correlation between test-set energy MAE and physical property prediction performance can be observed among energy-conserving models. All models are trained on MPTrj.
@@
Building on these insights, we present a novel MLIP called eSEN and training approach that achieves state-of-the-art (SOTA) performance on complex property prediction tasks. Specifically, our model is capable of running energy-conserving MD simulations for out-of-distribution systems (Figure 1 (a)). For materials stability prediction, eSEN achieves a leading F-1 score of $0.831$ and a $\kappa_{\mathrm{SRME}}$ of $0.321$ on the Matbench-Discovery benchmark (Riebesell et al., 2023; Póta et al., 2024). Previous models are only able to excel in one of these metrics ( Figure 1 (b,c)). On the MDR Phonon benchmark (Loew et al., 2024), SOTA results are found (Figure 1 (d)). Finally, eSEN achieves the highest test accuracy on the SPICE-MACE-OFF dataset (Kovács et al., 2023).
@@ 5.1 Direct-force prediction
Direct-force pre-training. Although direct-force models are not suitable for certain physical property prediction tasks, they may still offer advantages (Bigi et al., 2024; Amin et al., 2025). We demonstrate their efficiency can offer significant benefit as a pre-training strategy for a conservative model. Figure 3 shows the validation loss of 2-layer eSEN models trained on the MPTrj dataset: direct-force, conservative, and conservative fine-tuning from a pre-trained direct-force backbone. For conservative fine-tuning, we start from a direct-force model trained for 60 epochs, remove its direct-force prediction head, and fine-tune using conservative force prediction. The conservative fine-tuned model achieves a lower validation loss after being trained for 40 epochs compared to the from-scratch conservative model being trained for 100 epochs. The fine-tuning strategy also reduces the wallclock time for model training by 40%.
@@ 5.3 Smoothly varying PES
Table 2: Matbench-Discovery benchmark results of compliant models (trained only on MPtrj or its subset) with results on the unique prototype split. MAE is in units of eV/atom. ($\uparrow$) stands for higher the better. ($\downarrow$) stands for lower the better.
Model eSEN-30M eqV2 S DeNS Orb MPtrj SevenNet-l3i5 SevenNet-0 GRACE-2L (r6) MACE-MP-0 CHGNet M3GNet F1 $\uparrow$ 0.831 0.815 0.765 0.760 0.724 0.691 0.669 0.613 0.569 DAF $\uparrow$ 5.260 5.042 4.702 4.629 4.252 4.163 3.777 3.361 2.882 Precision $\uparrow$ 0.804 0.771 0.719 0.708 0.650 0.636 0.577 0.514 0.441 Recall $\uparrow$ 0.861 0.864 0.817 0.821 0.818 0.757 0.796 0.758 0.803 Accuracy $\uparrow$ 0.946 0.941 0.922 0.920 0.904 0.896 0.878 0.851 0.813 MAE $\downarrow$ 0.033 0.036 0.045 0.044 0.048 0.052 0.057 0.063 0.075 R2 $\uparrow$ 0.822 0.788 0.756 0.776 0.750 0.741 0.697 0.689 0.585 $\kappa_{\mathrm{SRME}}$downarrow 0.321 1.665 1.725 0.550 0.767 0.525 0.647 1.717 1.412
@@ 6.1 Matbench Discovery
The Matbench-Discovery benchmark evaluates a model’s ability to predict ground-state (0 K) thermodynamic stability through geometry optimization and energy prediction. It is a widely used benchmark for evaluating ML models in materials discovery. The compliant benchmark only includes models trained on the MPTrj (Jain et al., 2013; Deng et al., 2023) dataset or its subset, which facilitate a fair comparison of model architectures. The F1 score is the primary metric used to rank models. We train an eSEN with 30M parameters on MPTrj for 60 epochs of direct-force pre-training and 40 epochs of conservative fine-tuning. DeNS (Liao et al., 2024) is used during direct-force pre-training. As shown in Table 2, eSEN-30M achieves an F1 score of $0.831$—the highest among all compliant models.
@@
As shown in Table 2, our model achieves a $\kappa_{\mathrm{SRME}}$ of $0.321$ under the default evaluation protocol proposed by Póta et al. 2024. Notably, our model excels in both the F1 score and $\kappa_{\mathrm{SRME}}$, while all previous models only achieve SOTA performance on one or the other of these metrics.
@@ 6.2 MDR phonon benchmark
The MDR Phonon benchmark (Loew et al., 2024) assesses the performance of MLIPs in predicting key phonon properties, including maximum phonon frequency ($\omega_{\text{max}}$), entropy ($S$), free energy ($F$) and heat capacity at constant volume ($C_{V}$). The evaluation follows the testing protocol outlined by Loew et al. 2024. Table 3 shows the resulting MAE of our model and those of several other models [^2] In addition to our model, we also run the evaluation for GRACE-2L (r6) (Bochkarev et al., 2024), SevenNet-l3i5 (Park et al., 2024), Orb MPTrj (Neumann et al., 2024), and eqV2 S DeNS (Liao et al., 2023; Barroso-Luque et al., 2024), which were not included in the work by Loew et al. 2024.. eSEN achieves SOTA results, demonstrating a significant improvement over SevenNet, the second most accurate model.
@@
Table 3: Summary of the MAE for the MPTrj-trained models on the MDR Phonon benchmark. Metrics include maximum phonon frequency (MAE($\omega_{\mathrm{max}}$), in K), the vibrational entropy (MAE($S$), in J/K/mol), the Helmholtz free energy (MAE($F$), in kJ/mol), and the heat capacity at constant volume (MAE($C_{V}$), in J/K/mol).
Model MAE($\omega_{\text{max}}$) MAE($S$) MAE($F$) MAE($C_{V}$) M3GNet 98 150 56 22 CHGNet 89 114 45 21 MACE 61 60 24 13 GRACE-2L (r6) 40 25 9 5 SevenNet-0 40 48 19 9 SevenNet-l3i5 26 28 10 5 eSEN-30M 21 13 5 4 Direct-force models Orb MPTrj [0.01 Å] 309 476 64 181 Orb MPTrj [0.2 Å] 61 34 11 8 eqV2 S DeNS [0.01 Å] 280 224 54 94 eqV2 S DeNS [0.2 Å] 58 26 8 8
In physical phonon calculations, we expect the results to converge as the displacement goes to zero. By examining the resulting phonon band structure, we can gain insight into this behavior. Figure 5 presents the predicted phonon band structure and density of states for three representative materials using eSEN. The predicted phonon bands exhibit convergence as the displacement decreases. In contrast, Figures Figure C.11 and Figure C.12 display the phonon bands for the same three materials predicted using eqV2 S DeNS (direct-forces), which not only fail to demonstrate convergence but also exhibit significant errors, including missing acoustic branches and spurious imaginary frequencies.
@@ 7 Related works
MLIP architectures have made significant progress since their initial proposal (Behler and Parrinello, 2007). These architectures are usually symmetry-preserving (Smith et al., 2017; Schütt et al., 2017; Gilmer et al., 2017; Chmiela et al., 2017; Artrith et al., 2017; Unke and Meuwly, 2018; Zhang et al., 2018; Zubatyuk et al., 2019; Smith et al., 2020; Kovács et al., 2021), with increasingly expressive atom environment embeddings and message-passing operations (Gasteiger et al., 2020b, 2021; Schütt et al., 2021; Liu et al., 2021; Unke et al., 2021a; Chen and Ong, 2022; Deng et al., 2023; Cheng, 2024). Notably, equivariant architectures based on spherical harmonics representations (Thomas et al., 2018; Thölke and De Fabritiis, 2021; Batzner et al., 2022; Musaelian et al., 2022; Batatia et al., 2022; Passaro and Zitnick, 2023; Liao et al., 2023; Bochkarev et al., 2024; Park et al., 2024; Batatia et al., 2025) have shown strong performance on large-scale datasets. Meanwhile, the high computational cost of these architectures has sparked significant interest in scalable architectures that may not respect physical principles such as energy conservation (Langer et al., 2024; Brehmer et al., 2024; Hu et al., 2021; Yang et al., 2024; Qu and Krishnapriyan, 2024; Neumann et al., 2024). These models have demonstrated strong performance in accuracy, scalability, and relaxation tasks (Chanussot et al., 2021; Riebesell et al., 2023). While their non-physical nature may make them unsuitable for direct usage in some physical property prediction tasks, they may still provide benefit by using the pre-training strategy proposed in this paper, distilling them to conservative models (Amin et al., 2025), or combining them with a conservative model in a multiple-time-step integration scheme (Bigi et al., 2024).
@@ A.2 Phonon calculation protocols
Harmonic and anharmonic phonon calculations and solutions to the Wigner transport equation (Simoncelli et al., 2022) used for the thermal conductivity benchmark ($\kappa_{\text{SRME}}$) values given in Table 2 were carried out using the supercell method with finite differences implemented in Phono3py (Togo et al., 2015, 2023). The calculations followed the protocol described in the Matbench-Discovery benchmark (Riebesell et al., 2023; Póta et al., 2024). For eSEN-30M, an even lower $\kappa_{\mathrm{SRME}}$ of $0.298$ is obtained when we adjust the evaluation parameter atom displacement from $0.03$ to $0.05$.
@@ B.2 Displacement values and their relation to phonon band structure predictions
Figure B.9 presents the MAEs on phonon calculations for eSEN, MACE, and eqV2 S DeNS as a function of increasing displacement values. As anticipated, both eSEN and MACE exhibit constant or slightly increasing MAE with respect to displacement. In contrast, eqV2 S DeNS displays a notable decrease in MAE with increasing displacement. Notably, when using a displacement of 0.2 Å, the resulting phonon benchmark MAE values for eqV2 S DeNS become comparable to those of conservative force models (Table 3).
While the prediction accuracy of thermodynamic properties such as maximum frequency, entropy, free energy, and heat capacity improves with increasing displacement for direct-force models, this improvement is deceptive and does not translate to accurate predictions of the underlying phonon band structure and density of states (DOS). As illustrated in Figure C.11, the predicted phonon bands and DOS for three selected materials exhibit significant errors, particularly in capturing the correct dispersion relations. Moreover, imaginary frequencies are commonly predicted at small displacement values, suggesting a rough energy landscape (i.e. the learned PES is not truly convex when it’s very close to the minima). The eqV2 S DeNS model also fails to accurately capture acoustic modes—those that go to zero linearly at the $\gamma$ point—which is due to a non-zero net force on the structure. In contrast, non-zero net force at energy local minima does not occur for conservative models by definition (Figure 5).
@@
The apparent paradox of direct-force models like eqV2 S DeNS failing to accurately capture phonon band structures while still achieving competitive accuracy for thermodynamic properties such as entropy, free energy, and heat capacity can be resolved by examining the underlying calculation methodology. These properties are computed using weighted integrals of the DOS. Additionally the metrics in the MDR are performed at room temperature (300 K). The predicted DOS in Figure C.11 and Figure C.12 at larger displacement values adequately captures the overall features of the DFT DOS, but does not reproduce finer details such as high density areas and fluctuations. This level of agreement is sufficient for accurate predictions because the Boltzmann-weighted integrals used in calculating thermal properties help to mitigate the impact of point-wise errors, making it less crucial to precisely capture fine details in the band structure and DOS Ackland et al. (1997); Van De Walle and Ceder (2002). Moreover, since properties are estimated at 300 K, models can achieve accurate predictions by prioritizing prediction accuracy of lower-frequency modes, which are more relevant for thermal property calculations, rather than attempting to capture higher-frequency modes. Although eqV2 S DeNS without a net-zero force contraint may not accurately capture acoustic phonon branches, its ability to predict vibrational thermodynamics at room temperature remains unaffected due to the relatively small number of acoustic phonon states at low frequencies.
As a comparison with eqV2 S DeNS, Figures Figure C.13 and Figure C.14 shows the predicted phonon dispersion and DOS for the same three materials using eSEN with direct-force prediction. Although the results still exhibit some of the characteristic artifacts of direct-force models, such as convergence at larger displacements and the absence of acoustic modes, these issues are less pronounced compared to eqV2 S DeNS. Moreover, the predicted phonon bands and DOS are significantly improved, providing a more accurate representation of the DFT reference values. The better approximation of phonon bands and a lower tendency to predict imaginary frequencies highlight the importance of a smoothly-varying model, even without being conservative.

### v2

@@ Abstract
Machine learning interatomic potentials (MLIPs) have become increasingly effective at approximating quantum mechanical calculations at a fraction of the computational cost. However, lower errors on held out test sets do not always translate to improved results on downstream physical property prediction tasks. We propose testing MLIPs on their practical ability to conserve energy during molecular dynamic simulations. If passed, improved correlations are found between test errors and their performance on physical property prediction tasks. We identify choices which may lead to models failing this test, and use these observations to improve upon highly-expressive models. The resulting model, eSEN, provides state-of-the-art results on a range of physical property prediction tasks, including materials stability prediction, thermal conductivity prediction, and phonon calculations.
@@ 1 Introduction
Figure 1: (a) Energy conservation in MD simulations. Direct-force models (Orb, eqV2) and CHGNet fail to conserve. (b) A higher F1 score on the Matbench-Discovery strongly correlates with a lower test-set energy MAE. (c) Test-set energy MAE and $\kappa_{\mathrm{SRME}}$ on the Matbench-Discovery benchmark. (d) Test-set energy MAE and vibrational entropy MAE on the MDR Phonon benchmark. Our model (eSEN) achieves the best performance on all benchmarks. A higher correlation between test-set energy MAE and physical property prediction performance can be observed among energy-conserving models. All models are trained on MPTrj.
@@
Building on these insights, we present a novel MLIP called eSEN and training approach that achieves state-of-the-art (SOTA) performance on complex property prediction tasks. Specifically, our model is capable of running energy-conserving MD simulations for out-of-distribution systems (Figure 1 (a)). For materials stability prediction, eSEN achieves a leading F1 score of $0.831$ and a $\kappa_{\mathrm{SRME}}$ of $0.340$ on the compliant Matbench-Discovery benchmark (Riebesell et al., 2023; Póta et al., 2024). Previous models are only able to excel in one of these metrics ( Figure 1 (b,c)). We also achieve a SOTA F1 score of $0.925$ and $\kappa_{\mathrm{SRME}}$ of $0.170$ on the non-compliant category. On the MDR Phonon benchmark (Loew et al., 2024), SOTA results are found (Figure 1 (d)). Finally, eSEN achieves the highest test accuracy on the SPICE-MACE-OFF dataset (Kovács et al., 2023).
@@ 5.1 Direct-force prediction
Direct-force pre-training. Although direct-force models are not suitable for certain physical property prediction tasks, they may still offer advantages (Bigi et al., 2024; Amin et al., 2025). We demonstrate their efficiency can offer significant benefit as a pre-training strategy for a conservative model. Figure 3 shows the validation loss of 2-layer eSEN models trained on the MPTrj dataset: direct-force, conservative, and conservative fine-tuning from a pre-trained direct-force backbone. We start from a direct-force model trained for 60 epochs, remove its direct-force prediction head, and fine-tune using conservative force prediction. The conservative fine-tuned model achieves a lower validation loss after being trained for 40 epochs compared to the from-scratch conservative model being trained for 100 epochs. The fine-tuning strategy also reduces the wallclock time for model training by 40%. The strategy of combining direct-force pre-training and conservative fine-tuning is also shown to be effective under different data/model settings (Bigi et al., 2024).
@@ 5.3 Smoothly varying PES
Table 2: Matbench-Discovery benchmark results of compliant models (trained only on MPtrj or its subset) with results on the unique prototype split. MAE is in units of eV/atom. ($\uparrow$/$\downarrow$) stands for higher/lower the better.
Metric eSEN-30M-MP eqV2 S DeNS MatRIS-MP AlphaNet-MP DPA3-v2-MP ORB v2 MPtrj SevenNet-l3i5 GRACE-2L-MPtrj MACE-MP-0 CHGNet M3GNet F1 $\uparrow$ 0.831 0.815 0.809 0.799 0.786 0.765 0.760 0.691 0.669 0.613 0.569 DAF $\uparrow$ 5.260 5.042 5.049 4.863 4.822 4.702 4.629 4.163 3.777 3.361 2.882 Precision $\uparrow$ 0.804 0.771 0.772 0.743 0.737 0.719 0.708 0.636 0.577 0.514 0.441 Accuracy $\uparrow$ 0.946 0.941 0.938 0.933 0.929 0.922 0.920 0.896 0.878 0.851 0.813 MAE $\downarrow$ 0.033 0.036 0.037 0.041 0.039 0.045 0.044 0.052 0.057 0.063 0.075 R2 $\uparrow$ 0.822 0.788 0.803 0.745 0.804 0.756 0.776 0.741 0.697 0.689 0.585 $\kappa_{\mathrm{SRME}}$downarrow 0.340 1.676 0.861 1.31 0.959 1.725 0.550 0.525 0.647 1.717 1.412 RMSD $\downarrow$ 0.0752 0.0757 0.0773 0.1067 0.0823 0.1007 0.0847 0.0897 0.0915 0.0949 0.1117
Table 3: Matbench-Discovery benchmark results of non-compliant models with results on the unique prototype split.
Model eSEN-30M-OAM eqV2-M-OAM ORB v3 SevenNet-MF-ompa DPA3-v2-OpenLAM GRACE-2L-OAM MatterSim-v1-5M MACE-MPA-0 GNoME F1 $\uparrow$ 0.925 0.917 0.905 0.901 0.890 0.880 0.862 0.852 0.829 DAF $\uparrow$ 6.069 6.047 5.912 5.825 5.747 5.774 5.852 5.582 5.523 Precision $\uparrow$ 0.928 0.924 0.904 0.879 0.879 0.883 0.895 0.853 0.844 Accuracy $\uparrow$ 0.977 0.975 0.971 0.969 0.966 0.963 0.959 0.954 0.955 MAE $\downarrow$ 0.018 0.020 0.024 0.021 0.022 0.023 0.024 0.028 0.035 R2 $\uparrow$ 0.866 0.848 0.821 0.867 0.869 0.862 0.863 0.842 0.785 $\kappa_{\mathrm{SRME}}$downarrow 0.170 1.771 0.210 0.317 0.687 0.294 0.574 0.412 N/A RMSD $\downarrow$ 0.0608 0.0691 0.0750 0.0639 0.0679 0.0666 0.0733 0.0731 N/A
@@ 6.1 Matbench Discovery
The Matbench-Discovery benchmark evaluates a model’s ability to predict ground-state (0 K) thermodynamic stability through geometry optimization and energy prediction. It is a widely used benchmark for evaluating ML models in materials discovery. The compliant benchmark only includes models trained on the MPTrj (Jain et al., 2013; Deng et al., 2023) dataset or its subset, which facilitate a fair comparison of model architectures. The F1 score is the primary metric used to rank models. We train an eSEN with 30M parameters on MPTrj for 60 epochs of direct-force pre-training and 40 epochs of conservative fine-tuning. DeNS (Liao et al., 2024) is used during direct-force pre-training. As shown in Table 2, eSEN-30M-MP achieves an F1 score of $0.831$—the highest among all compliant models. eSEN-30M-MP also achieves the lowest root mean square deviation (RMSD) when comparing the relaxed structures to the ground truth DFT reference.
@@
As shown in Table 2, our model achieves a $\kappa_{\mathrm{SRME}}$ of $0.340$ under the default evaluation protocol proposed by Póta et al. 2024. Notably, our model excels in both the F1 score and $\kappa_{\mathrm{SRME}}$, while all previous models only achieve SOTA performance on one or the other of these metrics.
The non-compliant Matbench-Discovery benchmark includes models trained on datasets other than MPTrj. eSEN-30M-OAM is an eSEN model with 30 million parameters pre-trained on the OMat24 (Barroso-Luque et al., 2024) dataset then fine-tuned on the subsampled Alexandria (sAlex) dataset (Barroso-Luque et al., 2024; Schmidt et al., 2024) and MPTrj dataset. As shown in Table 3, eSEN-30M-OAM achieves the best performance among all non-compliant models with an F1 score of $0.925$, a $\kappa_{\mathrm{SRME}}$ of $0.170$, and an RMSD of $0.0608$, significantly advancing state-of-the-art.
@@ 6.2 MDR phonon benchmark
The MDR Phonon benchmark (Loew et al., 2024) assesses the performance of MLIPs in predicting key phonon properties, including maximum phonon frequency ($\omega_{\text{max}}$), entropy ($S$), free energy ($F$) and heat capacity at constant volume ($C_{V}$), for around 10,000 materials. The evaluation follows the testing protocol outlined by Loew et al. 2024. Table 4 shows the resulting MAE of our model and those of several other models [^2] In addition to our model, we also run the evaluation for GRACE (Bochkarev et al., 2024), SevenNet (Park et al., 2024), Orb (Neumann et al., 2024), and eqV2-S-DeNS (Liao et al., 2023; Barroso-Luque et al., 2024), which were not included in the work by Loew et al. 2024.. eSEN achieves SOTA results in both compliant and non-compliant categories.
@@
Table 4: Summary of model performance on the MDR Phonon benchmark. Metrics include maximum phonon frequency (MAE($\omega_{\mathrm{max}}$), in K), the vibrational entropy (MAE($S$), in J/K/mol), the Helmholtz free energy (MAE($F$), in kJ/mol), and the heat capacity at constant volume (MAE($C_{V}$), in J/K/mol). *The OMat24 dataset uses a slightly different DFT setting from the DFT setting of the MDR Phonon benchmark.
Compliant models MAE($\omega_{\text{max}}$) MAE($S$) MAE($F$) MAE($C_{V}$) M3GNet 98 150 56 22 CHGNet 89 114 45 21 MACE 61 60 24 13 GRACE-2L (r6) 40 25 9 5 SevenNet-0 40 48 19 9 SevenNet-l3i5 26 28 10 5 eSEN-30M-MP 21 13 5 4 Direct-force models Orb MPTrj [0.01 Å] 309 476 64 181 Orb MPTrj [0.2 Å] 61 34 11 8 eqV2-S-DeNS [0.01 Å] 280 224 54 94 eqV2-S-DeNS [0.2 Å] 58 26 8 8 Non-compliant models eqV2-M-OAM [0.01 Å] 780 403 241 100 eqV2-M-OAM [0.2 Å] 50 25 7 9 MatterSim 17 15 5 3 GRACE-2L-OAM 19 14 5 4 SevenNet-MF-ompa 15 8 3 3 eSEN-30M-OAM 15 10 4 3
In physical phonon calculations, we expect the results to converge as the displacement goes to zero. By examining the resulting phonon band structure, we can gain insight into this behavior. Figure 5 presents the predicted phonon band structure and density of states for three representative materials using eSEN. The predicted phonon bands exhibit convergence as the displacement decreases. In contrast, Figures Figure C.11 and Figure C.12 display the phonon bands for the same three materials predicted using eqV2-S-DeNS (direct-forces), which not only fail to demonstrate convergence but also exhibit significant errors, including missing acoustic branches and spurious imaginary frequencies.
While we find that the OMat-trained models (without sAlex/MPTrj finetuning) may provide a lower error on phonon prediction (7/7/2/2 for $\omega_{\mathrm{max}}$/$S$/$F$/$C_{V}$ MAE), we refrain from direct comparison due to mismatch in level of theory. We attribute this result to the softening issue of the sAlex and MPTrj dataset (Deng et al., 2025; Barroso-Luque et al., 2024), which OMat24 addresses. We refer interested readers to Deng et al. 2025 and Barroso-Luque et al. 2024 for a detailed discussion on the softening issue of some DFT datasets and its relation to phonon properties.
@@ 7 Related works
MLIP architectures have made significant progress since their initial proposal (Behler and Parrinello, 2007). These architectures are usually symmetry-preserving (Smith et al., 2017; Schütt et al., 2017; Gilmer et al., 2017; Chmiela et al., 2017; Artrith et al., 2017; Unke and Meuwly, 2018; Zhang et al., 2018; Zubatyuk et al., 2019; Smith et al., 2020; Kovács et al., 2021), with increasingly expressive atom environment embeddings and message-passing operations (Gasteiger et al., 2020b, 2021; Schütt et al., 2021; Liu et al., 2021; Unke et al., 2021a; Chen and Ong, 2022; Deng et al., 2023; Cheng, 2024; Yin et al., 2025). Notably, equivariant architectures based on spherical harmonics representations (Thomas et al., 2018; Thölke and De Fabritiis, 2021; Batzner et al., 2022; Musaelian et al., 2022; Batatia et al., 2022; Passaro and Zitnick, 2023; Liao et al., 2023; Bochkarev et al., 2024; Park et al., 2024; Batatia et al., 2025) have shown strong performance on large-scale datasets. Meanwhile, the high computational cost of these architectures has sparked significant interest in scalable architectures that may not respect physical principles such as energy conservation (Langer et al., 2024; Brehmer et al., 2024; Hu et al., 2021; Yang et al., 2024; Qu and Krishnapriyan, 2024; Neumann et al., 2024; Rhodes et al., 2025). These models have demonstrated strong performance in accuracy, scalability, and relaxation tasks (Chanussot et al., 2021; Riebesell et al., 2023). While their non-physical nature may make them unsuitable for direct usage in some physical property prediction tasks, they may still provide benefit by using the pre-training strategy proposed in this paper and Bigi et al. 2024, distilling them to conservative models (Amin et al., 2025), or combining them with a conservative model using multiple-time-step integration (Bigi et al., 2024).
@@ A.2 Phonon calculation protocols
Harmonic and anharmonic phonon calculations and solutions to the Wigner transport equation (Simoncelli et al., 2022) used for the thermal conductivity benchmark ($\kappa_{\text{SRME}}$) values given in Table 2 were carried out using the supercell method with finite differences implemented in Phono3py (Togo et al., 2015, 2023). The calculations followed the protocol described in the Matbench-Discovery benchmark (Riebesell et al., 2023; Póta et al., 2024). For eSEN-30M-MP, an even lower $\kappa_{\mathrm{SRME}}$ of $0.298$ is obtained when we adjust the evaluation parameter atom displacement from $0.03$ to $0.05$.
@@ B.2 Displacement values and their relation to phonon band structure predictions
Figure B.9 presents the MAEs on phonon calculations for eSEN, MACE, and eqV2-S-DeNS as a function of increasing displacement values. As anticipated, both eSEN and MACE exhibit constant or slightly increasing MAE with respect to displacement. In contrast, eqV2-S-DeNS displays a notable decrease in MAE with increasing displacement. Notably, when using a displacement of 0.2 Å, the resulting phonon benchmark MAE values for eqV2-S-DeNS become comparable to those of conservative force models (Table 4).
While the prediction accuracy of thermodynamic properties such as maximum frequency, entropy, free energy, and heat capacity improves with increasing displacement for direct-force models, this improvement is deceptive and does not translate to accurate predictions of the underlying phonon band structure and density of states (DOS). As illustrated in Figure C.11, the predicted phonon bands and DOS for three selected materials exhibit significant errors, particularly in capturing the correct dispersion relations. Moreover, imaginary frequencies are commonly predicted at small displacement values, suggesting a rough energy landscape (i.e. the learned PES is not truly convex when it’s very close to the minima). The eqV2-S-DeNS model also fails to accurately capture acoustic modes—those that go to zero linearly at the $\gamma$ point—which is due to a non-zero net force on the structure. In contrast, non-zero net force at energy local minima does not occur for conservative models by definition (Figure 5).
@@
The apparent paradox of direct-force models like eqV2-S-DeNS failing to accurately capture phonon band structures while still achieving competitive accuracy for thermodynamic properties such as entropy, free energy, and heat capacity can be resolved by examining the underlying calculation methodology. These properties are computed using weighted integrals of the DOS. Additionally the metrics in the MDR are performed at room temperature (300 K). The predicted DOS in Figure C.11 and Figure C.12 at larger displacement values adequately captures the overall features of the DFT DOS, but does not reproduce finer details such as high density areas and fluctuations. This level of agreement is sufficient for accurate predictions because the Boltzmann-weighted integrals used in calculating thermal properties help to mitigate the impact of point-wise errors, making it less crucial to precisely capture fine details in the band structure and DOS Ackland et al. (1997); Van De Walle and Ceder (2002). Moreover, since properties are estimated at 300 K, models can achieve accurate predictions by prioritizing prediction accuracy of lower-frequency modes, which are more relevant for thermal property calculations, rather than attempting to capture higher-frequency modes. Although eqV2-S-DeNS without a net-zero force contraint may not accurately capture acoustic phonon branches, its ability to predict vibrational thermodynamics at room temperature remains unaffected due to the relatively small number of acoustic phonon states at low frequencies.
As a comparison with eqV2-S-DeNS, Figures Figure C.13 and Figure C.14 shows the predicted phonon dispersion and DOS for the same three materials using eSEN with direct-force prediction. Although the results still exhibit some of the characteristic artifacts of direct-force models, such as convergence at larger displacements and the absence of acoustic modes, these issues are less pronounced compared to eqV2-S-DeNS. Moreover, the predicted phonon bands and DOS are significantly improved, providing a more accurate representation of the DFT reference values. The better approximation of phonon bands and a lower tendency to predict imaginary frequencies highlight the importance of a smoothly-varying model, even without being conservative.